OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form, and find the center and radius of the circle. 3x2 + 3y2 − 7x = 0

hero (hero):

complete the square

hero (hero):

Thats your hint

OpenStudy (anonymous):

Yeah I did that, you get (x^2 -7/3x +49/36) + y^2 = 49/36

hero (hero):

You need to put it in standard form

OpenStudy (anonymous):

So you tell me what subtracts into -7/3x and multiplies into 49/36

OpenStudy (anonymous):

No I don't.. It's a circle.

hero (hero):

Yes, you do.

hero (hero):

You're almost done.

OpenStudy (anonymous):

Standard form of a circle is (x-h)^2 + (y-k)^2 = r^2

hero (hero):

It's not that difficult

hero (hero):

Exactly. You need to put the expressions with the x's in that same form

OpenStudy (anonymous):

Yeah, so can you tell me what subtracts into 7/3 and multiplies into 49/36. Because it's obviously not Radical 7/3

hero (hero):

Basically you're telling me you're having trouble finding the perfect square form of the expression.

OpenStudy (anonymous):

basically I am telling you I don't know the Binomial expression. I have the polynomial quadratic, I just can't fathom any number that completes that into a binomial expression.

hero (hero):

I can't fathom how you completed the square without figuring out what the binomial expression is. I'll finish it for you.

OpenStudy (anonymous):

nvm got it.

hero (hero):

Well, what is it then?

OpenStudy (anonymous):

(x-7/6)^2 + y^2 = 49/36 so center is 7/6,0 radius is 7/6

hero (hero):

That is correct.

hero (hero):

I wonder what it was that was holding you back.

OpenStudy (anonymous):

My brain fart probably.. and the fact I hate staring at fractions.

hero (hero):

If I were you, I'd write the final form this way: \[(x - \frac{7}{6})^2 + y^2 = (\frac{7}{6})^2\]

OpenStudy (anonymous):

The professor wants it in r^2 format so I keep it in r^2

hero (hero):

That is r^2 form

OpenStudy (anonymous):

No I mean he wants it completely written out.,

hero (hero):

It's not that big a deal anyway

hero (hero):

If I were the instructor, either way would be fine with me.

hero (hero):

Just as long as you know how to do it.

OpenStudy (anonymous):

Well you aren't a raging asian.. so I am not going to fight with him you know?

hero (hero):

Raging asian? I haven't seen one of those since a kid got smacked by the teacher.

OpenStudy (anonymous):

Thats what I feel he will do.. haha